Under the approximate model of a lead acid battery, what is the cell voltage of a lead acid battery if the concentration of sulphuric acid is 0.1M?
Answer #1

Solution:

This can be calculated using the simple Nernst equation given below.

Ecell = E0 - (0.591 V / n) log [H+]

Since the overall number of electron transfer involved in a lead-acid battery is 2 and the standard cell potential is 2 V, we can write the above equation as follows

Ecell = 2 V - (0.591 V / 2) log (0.1)

Ecell = 2 - (0.0296) (-1) V

Ecell = 2 - (-0.0296) V

Ecell = 2.0296 V